OpenStudy (anonymous):

find axis of symmetry and vertex y=3x^2-12x+1 y=-2x^2+7x-21

sam (.sam.):

y=3x^2-12x+1 y=3(x^2-4x)+1 y=3(x-2)^2-4(3)+1 y=3(x-2)^2-11 Vertex at (2,-11) Symmetry at x=2

sam (.sam.):

y=-2x^2+7x-21 y=-2(x^2-7x/2)-21 y=-2(x-7/4)^2-49/16 (-2)-21 y=-2(x-7/4)^2-119/8 Vertex at (7/4,-119/8) Symmetry at x=7/4

OpenStudy (anonymous):

hank you